# Stop! In the Name of Flaws: Disentangling Personal Names and Sociodemographic Attributes in NLP

### v1

@@ 4 Ethical Issues
Harms can be broadly described as a setback in the interests or progress of people due to, e.g., the outcomes of an automatic process (Feinberg, 1984). Group-level harms are experienced collectively by people in a sociodemographic group, while individual harms (which might result from group membership) are experienced at the person-person or person-technology level. Inferring gender from names frequently misgenders trans people and erases non-binary people (Keyes, 2018). This perpetrates group-level erasure, as well as individual harms including damaging autonomy and dignity (Mcnamarah, 2020), inflicting psychological harms (Dev et al., 2021), and a failure to show recognition respect to people (Darwall, 1977). Certain type of name-based classification (e.g., of persecuted ethnic or religious groups) can threaten individual safety, and when NLP infrastructure is used for surveillance and targeting, this also threatens the safety of entire groups of people (Wadhawan, 2022). NLP systems reinforce group-level structural discrimination in other ways as well; studies of racial disparities in academic publishing have been shown to systematically discount the intellectual contributions of Black researchers (Kozlowski et al., 2022).
@@
In their work on name-based gender classification, Van Buskirk et al. (2023) note that for names with no available data, assigning the majority class (in their case, male) maximizes accuracy, but results in 0% error for the male class and 100% error for any other classes. For non-binary people, who are generally excluded from gender classification by design, the error rate is also almost always 100%. As for name-based race/ethnicity classifiers, Lockhart et al. (2023) show that people who self-identify as Filipino, Black, or Middle Eastern and North African, are misrecognized 55-75% of the time, as compared to those who identify as white, Chinese, or Korean, who are mislabelled less than 10% of the time. As described above, misrecognition errors cause harms, which are then disproportionately experienced by these individuals. We echo the conclusions of Lockhart et al. (2023) and Mihaljević et al. (2019), i.e., that inclusive analyses are only possible when names are no longer used as a proxy to infer individuals’ gender or race/ethnicity.
@@
The erasure of identities and the flattening of variation in naming customs leads to representational harms, which include the reinforcement of essentialist categories and power structures (Chien and Danks, 2024). These harms primarily affect sociodemographic groups, e.g., non-binary people, who are often incorrectly and unjustly treated as a novel social phenomenon and further experience hermeneutical injustice, i.e., unfair obstacles in “making sense of their social experience” (Medina, 2017). Groups of people with a certain name are often subject to a different type of representational harm, i.e., stereotyping. For instance, the name Kevin is associated with lower socioeconomic class in Germany (Kaiser, 2010). This stereotype, if encoded in an NLP system, could even lead to quality-of-service differentials, as class is a sociodemographic characteristic that correlates with lower NLP performance in other contexts (Curry et al., 2024).
@@
Names are a site for enforcing institutional power, as seen in “real name” policies (Haimson and Hoffmann, 2016), the (non-consensual) permanence of names in data infrastructure including Google Scholar (Speer, 2021), governmental name regulation (Te Tari Taiwhenua, 2021), and the “collective delusion” of legal names, at least in the U.S. (Baker and Green, 2021). Names are also regulated socially through norms and expectations, many of which end up baked into our NLP systems. We exercise power as NLP researchers and practitioners via our assumptions, which may reify sociodemographic categories, codify (or dismantle) associations between names and these categories, and create infrastructure that harms people at scale through surveillance or mislabelling. Knowles et al. (2016) open-sourced their name-based gender inference tool, and Vogel and Jurafsky (2012) published (binary) gender labels with names of authors of NLP papers, which continue to be used in research (Mohammad, 2020). This data reflects folk assumptions about gender, i.e., that it is binary, immutable and in perfect correspondence with names Keyes (2018); Cao and Daumé III (2021). These datasets also deadname and misgender scientists from the NLP community, some of whom have spoken about its harms (Mielke, 2024). Transgender people can only be counted in such a system if they conform to normative expectations (Johnson, 2016; Konnelly, 2021) and if not, the burden is disproportionately on them to seek redress. Even Asr et al. (2021)—a system relying on name-based gender inference that considers gender beyond the binary and does not publicly misgender individuals—does not shift power, as workarounds are a patch rather than built-in to the method; gender inference still relies on APIs that use binary gender and mistakes (typically, famous non-binary people) are manually corrected. As all these examples show how power remains centralized, we echo previous calls to reimagine and reconfigure power relations in service of user autonomy (Keyes et al., 2019; Blodgett et al., 2020; Hanna and Park, 2020).
@@ 5 Guiding Questions and Recommendations
It is acceptable to investigate what concepts NLP models associate with names, e.g., Madeleine with kindness. It is even acceptable to demonstrate that NLP models associate Marius with the pronoun he or with being male, and that these associations mirror common human associations (Caliskan et al., 2017; Crabtree et al., 2023). It is marginally acceptable to associate names with sociodemographic characteristics with imaginary people, e.g., drawing insights about gender bias more broadly based on how NLP models handle synthetic names of people assumed to be exclusively female; while doing so does not compromise people’s autonomy and dignity, it does further entrench hegemonic folk theories of names and people’s identities, which has cultural harms. Finally, it is unacceptable to present results about real people based solely on the assumption that their names provide a reliable signal about their identities, e.g., NLP papers authored by people named Madeleine and Marius cannot on their own provide trustworthy insights into gender and racial representation in the field, unless those specific individuals are asked about their gender.
@@
Names are rich in diversity, length, training data frequency, tokenization, associations, and the strengths of these associations, some of which has already been explored in prior work in NLP (Shwartz et al., 2020; Wolfe and Caliskan, 2021; Sandoval et al., 2023). The scope of what counts as a “name” should be carefully defined, based on the study’s context or the context of deployment. Each aspect must be operationalized and measured differently, and construct validity should be carefully considered. Ensure that pre-processing choices are contextualized and do not distort results, that names are understood within context, and that error can be quantified robustly.
@@
People’s identity and perceptions of them can differ, and these shape their experiences in various ways. Therefore, it is first necessary to decide which aspects are relevant for a study. Attempting to infer someone’s identity using names is simply unacceptable due to the range of methodological and ethical concerns we list in this paper. We echo onomastic advice from nearly 40 years ago (Weitman, 1981), i.e., that “inferences from names must be to the givers of these names, not to their bearers. What is more, inferences must always be to sociological formations (such as social classes, ethnic groups, historical generations, and the like), not to individual name-givers.” In addition to studying formations of name-givers, it can also be acceptable to study perceptions of identity based on names. For instance, numerous sociology papers have investigated racial and ethnic perceptions, as well as occupational stereotypes, based on names (King et al., 2006; Gaddis, 2017a, b). Again, we emphasize that perceptions based on names are also highly contextual and non-universal, e.g., not all societies believe that people should have names associated with positive qualities (Hough, 2016).
@@
It is essential to understand the geographical, temporal, and cultural context of data with names, and document this information for datasets, e.g., with datasheets (Gebru et al., 2021). What is the geographic, temporal cultural and political context of the name data, name-bearers, models and sociodemographic categories you use? Who are the people who will be impacted by your work, and what is their context? What do you know about the naming practices in these contexts and the hetereogeneity in these practices? Is error quantifiable here? We posit that it is unacceptable to use names without deeply engaging with context in these senses, and stress that ascribing contemporary Western identity categories to historical peoples without acknowledging the difference in contexts is reductive.
@@
For questions about people beyond identity, e.g., questions about how certain names are perceived by people, NLP may not be the only or best method available. We warn against technical solutionism (Green, 2021); researchers should reflect on whether their questions could be approached with interviews, case studies, fourth-world paradigms, and so on (Cameron, 2004). Qualitative methods can provide deeper, richer evidence while respecting people’s autonomy, dignity and context. If your questions are instead about NLP systems, then processing names with them is certainly necessary, but we note that methodological pluralism and interdisciplinarity can enrich our practice as NLP researchers and practitioners regardless (Wahle et al., 2023).
@@
Our paper provides a starting point for harms that are relevant to the use of names and sociodemographic characteristics in NLP, and we encourage transparency about methodological and ethical problems (Bietti, 2019; Hao, 2019). It is unacceptable to sideline these problems in the name of “social good” (Green, 2019; Greene et al., 2019; Bennett and Keyes, 2020), and rather than treating entire segments of the world as limitations of or future work for your research, we encourage changing the methods themselves, as in Lauscher et al. (2022). We recommend firmly grounding work in the ethical principles of autonomy, justice, and beneficence for people (Floridi and Cowls, 2019), which we note are sadly under-represented in machine learning research (Birhane et al., 2022).
@@
Descriptions of social phenomena and normative behaviour (i.e., assumptions and assertions that create and reinforce norms) are often conflated in NLP (Vida et al., 2023). This is the subtle but significant difference between showing that sociodemographic name associations in language models mirror the judgements of some group of humans, versus stating that model associations should mirror the judgements of some group of humans. The latter “cannot avoid creating and reinforcing norms” (Talat et al., 2022). Therefore, researchers should clearly distinguish descriptive and normative behaviours in the design, execution, and presentation of their experiments (Vida et al., 2023). System designers do have to make decisions about how systems should behave, i.e., they need to choose to perpetuate harmful structures in service of usability or to impose their own values on users and stakeholders when they take an advocacy position. This is an ethical dilemma in design that participatory methods and feminist epistemologies are uniquely positioned to help with (Bardzell, 2010).
@@ 7 Conclusion
We present the field with an overview of names and naming as discussed in other disciplines. We lay out background on naming practices around the world and describe how these practices create issues of validity (e.g., selection bias and construct validity) and ethical concerns (e.g., harms, cultural insensitivity) that affect NLP uses of names and sociodemographic characteristics. Finally, we present a list of guiding questions and normative suggestions towards addressing these concerns in future work involving names in NLP.
@@ Acknowledgments
We thank Lucy Li for recommending literature about personal names, sociodemographic characteristics, and social perceptions. We are also grateful to members of the Critical Media Lab Basel, Switzerland, and the Interdisciplinary Institute for Societal Computing, Germany, for their feedback on the ideas in this paper.

### v2

@@ 4 Ethical Issues
Harms can be broadly described as a setback in the interests or progress of people due to, e.g., the outcomes of an automatic process (Feinberg, 1984). Group-level harms are experienced collectively by people in a sociodemographic group, while individual harms (which might result from group membership) are experienced at the person-person or person-technology level. Inferring gender from names frequently misgenders trans people and erases non-binary people (Keyes, 2018). This perpetrates group-level erasure, as well as individual harms including damaging autonomy and dignity (Mcnamarah, 2020), inflicting psychological harms (Dev et al., 2021), and a failure to show recognition respect to people (Darwall, 1977). Certain types of name-based classification (e.g., of persecuted ethnic or religious groups) can threaten individual safety, and when NLP infrastructure is used for surveillance and targeting, this also threatens the safety of entire groups of people (Wadhawan, 2022). [^3] Regulation efforts such as the AI Act (Commission, 2021) in the EU try to mitigate this, but this does not apply to authoritarian regimes’ use of such technology (Briglia, 2021). NLP systems reinforce group-level structural discrimination in other ways as well; name-based studies of racial disparities in academia have been shown to systematically discount the intellectual contributions of Black researchers (Kozlowski et al., 2022).
@@
In their work on name-based gender classification, Van Buskirk et al. (2023) note that for names with no available data, assigning the majority class (in their case, male) maximizes accuracy, but results in 0% error for the male class and 100% error for any other classes. For non-binary people, who are generally excluded from gender classification by design, the error rate is also almost always 100%. As for name-based race/ethnicity classifiers, Lockhart et al. (2023) show that people who self-identify as Filipino, Black, or Middle Eastern and North African, are misrecognized 55-75% of the time, as compared to those who identify as white, Chinese, or Korean, who are mislabelled less than 10% of the time. As described above, misrecognition errors cause harms, which are then disproportionately experienced by these individuals. We echo the conclusions of Mihaljević et al. (2019) and Lockhart et al. (2023), i.e., that inclusive analyses are only possible when names are no longer used as a proxy to infer individuals’ gender or race/ethnicity.
@@
The erasure of identities and the flattening of variation in naming customs leads to representational harms, which include the reinforcement of essentialist categories and power structures (Chien and Danks, 2024). These harms primarily affect sociodemographic groups, e.g., non-binary people, who are often incorrectly and unjustly treated as a novel social phenomenon. Groups of people with a certain name are often subject to a different type of representational harm, i.e., stereotyping. For instance, the name Kevin is associated with lower socioeconomic class in Germany (Kaiser, 2010). This stereotype, if encoded in an NLP system, could lead to quality-of-service differentials, as class is a sociodemographic characteristic that correlates with lower NLP performance in other contexts (Curry et al., 2024).
@@
Names are a site for enforcing institutional power, as seen in “real name” policies (Haimson and Hoffmann, 2016), the (non-consensual) permanence of names in data infrastructure including Google Scholar (Speer, 2021), governmental name regulation (Te Tari Taiwhenua, 2021), and the “collective delusion” of legal names, at least in the U.S. (Baker and Green, 2021). Names are also regulated socially through norms and expectations, many of which end up baked into our NLP systems. We exercise power as NLP researchers and practitioners via our assumptions, which may reify sociodemographic categories, codify (or dismantle) associations between names and these categories, and create infrastructure that harms people at scale through surveillance or mislabelling. Knowles et al. (2016) open-sourced their name-based gender inference tool, and Vogel and Jurafsky (2012) published (binary) gender labels with names of authors of NLP papers, which continue to be used in research (Mohammad, 2020; Van Buskirk et al., 2023). This data reflects folk assumptions about gender, i.e., that it is binary, immutable and in perfect correspondence with names Keyes (2018); Cao and Daumé III (2021). These datasets also deadname and misgender scientists from the NLP community, some of whom have spoken about its harms (Mielke, 2024). Transgender people can only be counted in such a system if they conform to normative expectations (Johnson, 2016; Konnelly, 2021), and if not, the burden is disproportionately on them to seek redress. Even Asr et al. (2021)—a system relying on name-based gender inference that considers gender beyond the binary and does not publicly misgender individuals—does not shift power, as workarounds are a patch rather than built-in to the method; gender inference still relies on APIs that use binary gender, and mistakes (typically, famous non-binary people) are manually corrected. As all these examples show how power remains centralized, we echo previous calls to reimagine and reconfigure power relations in service of user autonomy (Keyes et al., 2019; Blodgett et al., 2020; Hanna and Park, 2020).
@@ 5 Guiding Questions and Recommendations
It is acceptable to investigate what concepts NLP models associate with names, e.g., Madeleine with kindness. It is even acceptable to demonstrate that NLP models associate Marius with the pronoun he or with being male, and that these associations mirror common human associations (Caliskan et al., 2017; Crabtree et al., 2023). It is marginally acceptable to associate names with sociodemographic characteristics using imaginary people, e.g., drawing insights about gender bias more broadly based on how NLP models handle synthetic names of people assumed to be exclusively female; while doing so does not compromise people’s autonomy and dignity, it does further entrench hegemonic folk theories of names and people’s identities, which has cultural harms. Finally, it is unacceptable to present results about real people based solely on the assumption that their names provide a reliable signal about their identities, e.g., NLP papers authored by people named Madeleine and Marius cannot on their own provide trustworthy insights into gender and racial representation in the field, unless those specific individuals are asked about their gender.
@@
Names are rich objects of study with variation in form, length, training data frequency, tokenization, associations, the strengths of these associations, and more. [^4] Some of these aspects have already been explored in prior work in NLP (Shwartz et al., 2020; Wolfe and Caliskan, 2021; Sandoval et al., 2023). Once you have decided what aspects to study, they must be operationalized and measured carefully, with attention to the context of the study or eventual system deployment. This includes the scope of what counts as a “name.” For instance, considering the use of English common nouns as names (e.g., Cloud) is particularly important when working with data from or systems deployed in China, where this naming practice is common (Chan, 2016). Ensure that pre-processing choices are contextualized and do not distort results, that names are understood within context, and that error can be quantified robustly in the given context. Thus, when measuring training data frequency of names, counting Cloud tokens as names must consider when it is used as a name and when it is used simply as a noun. Error could be quantified through manual analysis on a subset of the data.
@@
People’s identity and perceptions of them can differ, and these shape their experiences in various ways. Therefore, it is first necessary to decide which aspects are relevant for a study. Attempting to infer someone’s identity using names is simply unacceptable due to the range of methodological and ethical concerns we list in this paper. We echo onomastic advice from nearly 40 years ago (Weitman, 1981), i.e., that “inferences from names must be to the givers of these names, not to their bearers. What is more, inferences must always be to sociological formations (such as social classes, ethnic groups, historical generations, and the like), not to individual name-givers.” In addition to studying formations of name-givers, it can also be acceptable to study perceptions of identity based on names. For instance, numerous sociology papers have investigated racial and ethnic perceptions, as well as occupational stereotypes, based on names (King et al., 2006; Gaddis, 2017a, b). Again, we emphasize that perceptions based on names are also highly contextual and non-universal.
@@
It is essential to understand the geographical, temporal, and cultural context of data with names, and document this information for datasets, e.g., with datasheets (Gebru et al., 2021). What is the geographic, temporal, cultural and political context of the name data, name-bearers, models and sociodemographic categories you use? Who are the people who will be impacted by your work, and what is their context? What do you know about the naming practices in these contexts and the hetereogeneity in these practices? Are you quantifying error with self-reported data? We posit that it is unacceptable to use names without deeply engaging with context in these senses, and stress that ascribing contemporary Western identity categories to historical peoples without acknowledging the difference in contexts is reductive.
@@
For questions about human identity and perception based on names, NLP may not be the only or best method available. We warn against technical solutionism (Green, 2021); researchers should reflect on whether their questions could be approached with interviews, case studies, fourth-world paradigms, and so on (Cameron, 2004). Qualitative methods can provide deeper, richer evidence while respecting people’s autonomy, dignity and context. If your questions are instead about NLP systems, then processing names with them is certainly necessary, but we note that methodological pluralism and interdisciplinarity can enrich our practice as NLP researchers and practitioners regardless (Wahle et al., 2023).
@@
Our paper provides a starting point for harms that are relevant to the use of names and sociodemographic characteristics in NLP, and we encourage transparency about methodological and ethical problems (Bietti, 2019; Hao, 2019). It is unacceptable to sideline these problems in the name of “social good” (Green, 2019; Greene et al., 2019; Bennett and Keyes, 2020), and rather than treating entire segments of the world as limitations of or future work for your research, we encourage changing the methods themselves, as Lauscher et al. (2022) do with neopronouns. We recommend firmly grounding work in the ethical principles of autonomy, justice, and beneficence for people (Floridi and Cowls, 2019), which we note are sadly under-represented in machine learning research (Birhane et al., 2022).
@@
Descriptions of social phenomena are often conflated with normative behaviour (i.e., assumptions and assertions that create and reinforce norms) in NLP (Vida et al., 2023). This is the subtle but significant difference between showing that sociodemographic name associations in language models mirror the judgements of some group of humans, versus stating that model associations should mirror the judgements of some group of humans. The latter “cannot avoid creating and reinforcing norms” (Talat et al., 2022). Therefore, researchers should clearly distinguish descriptive and normative behaviours in the design, execution, and presentation of their experiments (Vida et al., 2023). System designers do have to make decisions about how systems should behave, i.e., they need to choose to perpetuate harmful structures in service of usability or to impose their own values on users and stakeholders when they take an advocacy position. This is an ethical dilemma in design that participatory methods and feminist epistemologies are uniquely positioned to help with (Bardzell, 2010).
@@ 7 Conclusion
We present the field with an overview of names and naming as discussed in other disciplines. We lay out background on naming practices around the world and describe how these practices create issues of validity (e.g., selection bias and construct validity) and ethical concerns (e.g., harms, cultural insensitivity), that affect NLP uses of names and sociodemographic characteristics. Finally, we present a list of guiding questions and normative suggestions towards addressing these concerns in future work involving names in NLP.
@@ Acknowledgments
We thank Lucy Li for recommending literature about personal names, sociodemographic characteristics, and social perceptions. We are also grateful to our reviewers, and to members of the Critical Media Lab Basel, Switzerland, and the Interdisciplinary Institute for Societal Computing, Germany, for their feedback on the ideas in this paper.